A motorcycle and a car leave an intersection at the same time. The motorcycle heads north at an average speed of 20 miles per hour, while the car heads east at an average speed of 48 miles per hour. Find an expression for their distance apart in miles at the end of t hours.


Answer by dolly(163)

Speed of the motorcycle heading toward north = 20mi/hr
Time travelled = t hrs
so distance covered = speed x time
= 20 t miles
Speed of the motorcycle heading toward east = 48mi/hr
Time travelled = t hrs
so distance covered = speed x time
= 48 t miles
North and east are at right angles.
So applying Pythagorean theorem, the distance between them is given by,
distance = sqrt( (48t)^2 + (20t)^2 )
= sqrt ( 2304 t^2 + 400 t^2 )
= sqrt ( 2704 t^2 )
= 52 t
So at the end of t hrs they are 52 t miles apart.
good luck!!!